How much, on average, do you expect to gain by investing $4,120 in the following stock: To buy the stock, you have to pay a $120 non-refundable service charge. There is a 60% chance that the firm will be bought by a large computer firm next year. If that is the case, there is a 50% chance that the stock holders will be bought out at double the present stock price, and a 50% chance that the stock holders will be bought out at 150% of the present stock price. If the firm is not bought by the large company, there is a 50% chance that the firm’s stock value will stay the same, and a 50% chance that the firm will go broke. If the firm goes broke, its stock will be totally worthless. Group of answer choices
Answer #1

Total invested - 4120 + 120 = 4240

Average price after considering all the senarios as calculated below: 5150

Net gain = 4240 - 5150 = $910.

A G Stock price Pay 4120 120 Non refundable 60% Firm will be bought 40% chance it wont be bought 60% 50% chance price will be